# Explaining GPT-4’s Schema of Depression Using Machine Behavior Analysis

### v1

@@ Abstract
[…]
keywords:
Depression, Language Models, Machine Behavior, GPT-4
@@ 1 Introduction
Large Language Models (LLMs), such as OpenAI’s GPT-4 or Google’s Gemini are starting to be actively used by thousands of individuals for mental health support, although we have a limited understanding of their outcomes and limitations. As of January 2024, the ChatGPT store [1] reported tens of thousands of conversations involving GPT therapy bots [2], and approximately 1 in 25 chatbot users seek mental health support as one of their primary uses of LLMs [3]. A recent survey in the Journal of Medical Economics revealed that over 1 in 10 clinicians in the USA use ChatGPT, and roughly 50% are interested in continuing its use for data entry, medical scheduling, or research [4].
The potential benefits of LLMs in the mental health domain are numerous. Mental health care often begins with psychodiagnostic assessment, where employing LLMs could play a crucial role for more timely and effective screening and assessment [5]. Present-day screening tools have limited sensitivity, standardization [6], and cultural adaptability [7, 8], plus people report increased willingness to use online mental health technologies available to them [5]. Further, the ability to administer assessments earlier in the care cycle can address the fact that diagnosis often lags behind the onset of symptoms by over a decade [9, 10, 11]. Earlier detection and treatment could mitigate financial barriers that arise from treating more severe conditions [12, 13, 14]. However, similar to treating mental disorders with unevaluated drugs, unevaluated use of LLMs for mental health support comes with many risks [15, 16].
Figure 1: Overview of Machine Behavior Evaluation to uncover the structure of depression in GPT-4. (1) GPT-4 is prompted with step-by-step instructions on performing depression assessment on human-written essays. (2) GPT-4 carried out assessments on data from Gu et al., 2023 [17], containing depression essays written by individuals. The model estimated the severity scores for PHQ-9 symptoms explicitly mentioned in the essay, followed by scoring the remaining symptoms that were implicit in the essay. (3) GPT-4’s estimated scores and spans are used to infer its internal structure of depressive symptoms, which revealed how the web of symptoms was connected, and what language GPT-4 reliably used to estimate depression severity.
@@
This paper presents a framework for evaluating LLMs’ “representation” of mental health. To do this, we apply principles from measurement theory [18] to infer GPT-4’s schema of depression (how it appears to organize, interpret, and relate depressive symptoms). In the process, we quantitatively evaluate two fundamental questions: (a) how GPT-4 identifies and estimates depression symptoms from language, and (b) its latent structure between the symptoms used to form an overall assessment. We utilize this framework to study GPT-4’s structural representation of depression for several reasons. First, depression is one of the most prevalent mental health conditions, affecting an estimated 10–15% of the population throughout their lifetime [19, 20], and is the leading cause of disability worldwide [21]. Second, there exists a vast body of research on depression’s symptoms, structure, and their inter-relationships, providing a strong theoretical reference against which to compare [22, 23, 24]. Lastly, depression is a highly heterogeneous disorder comprised of cognitive, affective, behavioral, and somatic symptoms [25]; depression is also highly co-morbid with other conditions, including anxiety [26], substance use [27], somatic disorders [28] and other internalizing conditions [29]. Thus, understanding GPT-4’s schema of depression suggests generalizations to many aspects and forms of psychopathology.
@@
While previous work on how GPT-4’s assessments are formed is limited, recent studies have evaluated GPT-4’s ability to predict mental disorder [30]. Empirical approaches that seek to explain GPT-4’s behavior are often limited to performance benchmarks and comparative studies [31], which measure the agreement between model’s responses and those of humans [32, 33] or experts [34]. Such methods can reveal what the model can do but do not explain how it performs such tasks. The responsible use of LLMs, particularly in critical applications that involve human health and well-being, requires understanding how they perform tasks —– questions that have traditionally fallen within the purview of fields like psychology and cognitive science.
@@
Over the years, psychology has developed and continually refined tools and methods to measure, interpret, and understand the human mind using mathematical and empirical methods designed with the intent of explaining human mental functioning through behavioral observation [35]. Such methods have uncovered schemas — the internal concepts underlying a process — embedded within the neural system from observations of human behaviors [36, 37, 38]. Drawing parallels from cognitive and behavioral psychology, our experiments are designed to collect extensive observations of GPT-4’s ability to assess depression. We analyze its generated text (responses) and the input prompts (stimuli) to infer GPT-4’s schema of depression. This inferred schema is used to explain GPT-4’s behavior in assessing depression [39]. To the best of our knowledge, this is the first work to perform machine psychology [40] to explain GPT-4’s functioning. We employ a robust methodological framework that can be extended to other LLMs to understand how psychopathology is represented within them. This is depicted in Figure 1 and described extensively under Methods (§4).
@@
The unique contributions of this study center on explaining GPT-4’s behavior when identifying depression risk within open-ended essays. In a sample of 955 individuals, we employ measurement theory to infer GPT-4’s schema of depression by analyzing (a) the relationships among key depression symptoms and (b) the role of each symptom in determining the degree of depression within the language model. This inferred schema is compared with one derived from self-reported scores to quantitatively identify similarities and differences in the conceptualization of depression as a network of interconnected symptoms. We used multivariate linear regression to assess how GPT-4 integrates explicitly mentioned factors in essays to estimate risks for unmentioned factors. We then applied differential language analysis [41, DLA] to identify the linguistic cues associated with GPT-4’s identification of depressive risk factors. Lastly, we evaluated whether GPT-4 shows evidence that it uses the phrases that it cited as explanations for producing its depression scores.
@@ 2 Results
GPT-4 was used to analyze language samples from 955 participants, who described their feelings of depression, with the goal of estimating symptom-level scores based on the PHQ-9 [42] questionnaire. This process involved two key steps. In the first step, the model identified symptoms explicitly mentioned in the participants’ essays, citing the corresponding text spans and assigning severity scores to each explicitly mentioned symptom. In the second step, the model inferred any remaining PHQ-9 symptoms that were not directly mentioned in the essays but could be implied from the context, providing a free-form rationale for each inferred symptom, along with a severity estimate.
To validate these estimates, participants also completed the PHQ-9 questionnaire, which assessed their symptom severity over the previous two weeks. Additionally, a random subsample of 209 participant responses was manually annotated by two psychology experts using a similar annotation process to that given to GPT-4. The comparison between GPT-4’s estimates and the participants’ self-reported PHQ-9 scores was conducted on the full dataset, while the comparison between GPT-4 and expert annotations, as shown in Table 1, was based on the subset of expert-annotated data.
[…]
Experts - GPT Self-report - GPT Experts - Self-report
Symptom $r$ $r$ $r$
PHQ 9 Total .81 .70 .68
Anhedonia .65 .59 .56
Depressed .81 .69 .66
Sleep .69 .54 .46
Fatigue .68 .59 .53
Appetite .49 .43 .43
Worthlessness/Guilt .70 .50 .53
Concentration .62 .48 .45
Psychomotor .31 .23 .14
Suicidal Thoughts .63 .31 .30
Figure 2: Similarities and differences between GPT-4’s and self-report’s symptom structure as symptom covariances.
(a) Pearson correlation between symptoms estimated by GPT-4 (left) and self-reported by participants (right). The average correlations of each symptom with rest are presented as column vectors for both GPT-4 and self-report.
(b) Difference between GPT-4’s and self-reports’ Symptom-Symptom correlations. 95% confidence interval of the difference in correlations from bootstrapped resampling over 500 trials is enclosed within brackets.
@@
Table 1 shows the average Pearson correlation ($r$) between GPT-4 estimates and experts for each symptom, as well as the correlation between GPT-4 and self-reported scores, computed on the subset assessed by the experts ($N=209$). GPT-4 demonstrated high agreement with self-report ($r=0.70$) and experts-judgments (avg $r=0.81$) on total PHQ-9 ratings. Across all symptoms, the assessors of depression essays (i.e., GPT-4 and the experts) demonstrated higher agreement with each other than either did with self-reported PHQ-9 scores. The absolute difference between GPT-4’s agreements with self-report and experts (difference between first two columns) was less than $0.10$ for all somatic symptoms (fatigue, appetite, and psychomotor agitation/retardation), except for issues related to Sleep ($\Delta=0.15$). All cognitive and emotional symptoms except for anhedonia had a difference in correlation scores over 10 Pearson points. Highest differences were seen with suicidal ideation ($\Delta=0.32$), worthlessness/guilt ($\Delta=0.20$) and concentration difficulties ($\Delta=0.14$). GPT-4’s agreement with experts ($r_{exp1}=0.80,r_{exp2}=0.81$) on total PHQ-9 score was very similar to the agreement between both the experts ($r=0.83$).
@@
The symptom-symptom associations from GPT-4 estimates as depicted by the heatmap (left) in 2(a) had positive Pearson correlations across the board ($r=0.23-0.78$). This was in agreement with the symptom-symptom associations from self-reported (right) scores ($r=0.38-0.78$). The effect sizes of the symptom-symptom relationships from GPT-4 estimates were similar to self-report for all but two symptoms, namely, psychomotor agitation/retardation and suicidal ideation. 2(b) clearly shows this difference between symptom-symptom associations from GPT-4 and self-report is statistically significant (95% CI doesn’t include 0.0). GPT-4-estimated severities of psychomotor agitation/retardation had a stronger association with other symptoms than self-report. In contrast, suicidal ideation was less related to other symptoms for GPT-4 than in the self-report. This is evident from the significantly lower association of suicidal ideation with the rest of the symptoms than what was observed in self-reported scores.
Table 2: Symptom’s item location ranking according to IRT over GPT-4 and self-report scores for zero and non-zero decision boundary. A lower rank value corresponds to the symptom being best at distinguishing lower values of depression, while a higher rank indicates being better at distinguishing higher values of depression. Ranks are accompanied by a 95% confidence interval estimated using bootstrapped resampling across individuals (N=955, num. trials=500).
PHQ9 Item Item Location Rank
GPT-4 Self-Report Item location difference
Anhedonia 3 [1, 3] 3 [3, 3] 0
Depressed Mood 1 [1, 2] 2 [1, 2] -1
Sleep 4 [4, 5] 4 [4, 5] 0
Fatigue 2 [1, 3] 1 [1, 2] 1
Appetite 8 [7, 8] 7 [7, 7] 1
Worthlessness / Guilt 6 [6, 6] 5 [4, 6] 1
Difficulty concentrating 5[4, 5] 6[4, 6] -1
Psychomotor 7[7, 8] 8[8, 8] -1
Suicidal Ideation 9[9, 9] 9[9, 9] 0
@@
We utilize Item Response Theory (IRT) to conduct a Differential Item Functioning (DIF) analysis on two sets of scores — self-report and GPT-4 estimates. The DIF ranks highlight the depression severity ranges where each item most reliably identifies the presence of its symptom. [43, 44, 45]. A lower-ranked items are best at distinguishing lower latent scores, and higher-ranked items for higher latent scores. A positive difference in an item’s rank between GPT-4 and self-report means that language-based assessment of this item by GPT-4 is more informative than self-report scores to capture higher depression scores. On the other hand, a negative difference implies that the language-based score of this item has lesser discriminant power than the self-report over higher depression scores. Table 2 shows the DIF ranks for GPT-4 and self-report along with its 95% confidence interval estimated through bootstrapped resampling across individuals over 500 trials.
@@
Overall, the distinguishing characteristics of items were in agreement between the two, with suicidal ideation, appetite, and psychomotor agitation/retardation having higher ranks (7-9), and anhedonia, depressed mood, and fatigue having lower ranks (1-3). All items had a very small 95% confidence interval, showing that these rank estimates have high reliability. Suicidal ideation (rank=9), sleep (rank=4), and anhedonia (rank=3) were the three items with no difference in ranks between GPT-4 and self-report. The difference in ranks of other items doesn’t indicate that language-based severity scores of GPT-4 possessed a significantly different characteristic compared to self-report, i.e. there is a high similarity in the item functioning of self-report and GPT-4 for PHQ-9.
Table 3: Contrast of explicit and implicit symptom mentions and convergence. Pearson Correlation ($r$) between GPT-4 and self-reported scores for all symptoms under conditions where symptoms were identified to be explicit and implicit (N=955) by GPT-4. The bar chart in % Explicit column represents the rate of explicit mentions according to GPT-4. Higher correlations were observed consistently when symptoms were identified to be explicitly mentioned.
Symptom Explicit $r$ Explicit vs. % Implicit Implicit $r$
Anhedonia .71 53.72% .43
Depressed Mood .70 96.44% .44
Sleep .52 11.31% .43
Fatigue .58 25.13% .51
Appetite .47 5.03% .33
Worthlessness / Guilt .57 24.19% .45
Concentration .50 9.42% .45
Psychomotor .33 3.35% .24
Suicidal Ideation .75 5.55% .21
Mean .57 26.02% .39
Explicit & Implicit Symptoms: GPT-4 is more precise in estimating explicit symptoms than implicit ones.
@@
Convergence of GPT-4 with self-report in samples identified with explicit mentions of symptoms was higher than unmentioned samples across all symptoms with an average $\Delta$ of +0.18 Pearson points (Table 3). Particularly, GPT-4’s estimation was significantly more ($\Delta\geq 0.25$ Pearson points) in agreement with self-report under explicit mention than implicit for anhedonia, depressed mood, and suicidal ideation. Language describing these three symptoms led to more precise GPT-4 estimates than language describing other symptoms, despite their varying rates of explicit mentions ($range=5.55\%-96.44\%$). The cardinal symptoms, ie., depressed mood and anhedonia, had notably higher rates of explicit mentions in the essays ($96\%$ and $53\%$ respectively) than other symptoms.
Figure 3: Effect of explicit symptom scores on implicit symptom scores as estimated by GPT-4. Wider bands represent larger effect sizes ($\beta$ values, i.e., coefficients of multivariate linear regression from predicting estimated implicit symptom score using all explicit symptom scores while holding other implicit symptoms constant). The 90% confidence interval of the regression coefficients was estimated using bootstrapped resampling over 500 trials. Confidence intervals that included 0.0 were dropped. Explicit refers to the symptoms estimated by GPT-4 when it identified explicit mention in essays and Implicit refers to the symptom scores when they were unmentioned in essays. Dotted bands represent a negative association.
Figure 4: GPT-4’s Language of Depression: Words from the human-written essays that were associated with being identified as explicit symptomatic markers (left) and higher severity score (right) as estimated by GPT-4. Positive Cohen’s d values are concordant with GPT-4-identified explicit markers $(p<0.05)$. Higher Pearson $r$ maps to higher depression severity and words with significant correlations $(p<0.05)$ are boldfaced. Both tests were Bonferroni corrected.
Interplay between explicitly mentioned symptoms and inferred symptoms

### v2

# Explaining GPTs’ Schema of Depression: A Machine Behavior Analysis

@@ Abstract
Use of large language models such as ChatGPT (GPT-4/GPT-5) for mental health support has grown rapidly, emerging as a promising route to assess and help people with mood disorders, like depression. However, we have a limited understanding of these language models’ schema of mental disorders, that is, how it internally associates and interprets symptoms of such disorders. In this work, we leveraged contemporary measurement theory to decode how GPT-4 and GPT-5 interrelates depressive symptoms, providing an explanation of how LLMs apply what they learn and informing clinical applications. We found that GPT-4: (a) had strong convergent validity with standard instruments and expert judgments (r=.70.81)(r=.70-.81), and (b) behaviorally linked depression symptoms with each other (symptom inter-correlates r=.23.78r=.23-.78) in accordance with established literature on depression; however, it (c) underemphasized the relationship between suicidality and other symptoms while overemphasizing psychomotor symptoms; and (d) suggested novel hypotheses of symptom mechanisms, for instance, indicating that sleep and fatigue are broadly influenced by other depressive symptoms, while worthlessness/guilt is only tied to depressed mood. GPT-5 showed a slightly lower convergence with self-report, a difference our machine-behavior analysis makes interpretable through shifts in symptom–symptom relationships. These insights provide an empirical foundation for understanding language model’s mental health assessments and demonstrate a generalizable approach for explainability in other models and disorders. Our findings can guide key stakeholders to make informed decisions for effectively situating these technologies in the care system.
Keywords: Depression, Large Language Models, GPT, Machine Behavior
@@ 1 Introduction
Large Language Models (LLMs), such as OpenAI’s GPT-4 or Google’s Gemini are starting to be actively used by thousands of individuals [18] for mental health support, although we have a limited understanding of their outcomes and limitations. A recent survey in the Journal of Medical Economics revealed that over 1 in 10 clinicians in the USA use ChatGPT, and roughly 50% are interested in continuing its use for data entry, medical scheduling, or research [42]. However, similar to treating mental disorders with drugs that are not understood, there are many risks in using LLMs for mental health support without explanation of how they behave in the domain [45, 8].
This paper presents a framework for evaluating and explaining LLMs’ “representation” of mental health. To do this, we apply principles from measurement theory [37] to infer GPT-4’s and GPT-5’s schema of depression (how it appears to organize, interpret, and relate depressive symptoms) based on its behavior. In the process, we quantitatively evaluate two fundamental questions: (a) how does GPT identify and estimate depression symptoms from language, and (b) what is its latent structure between the symptoms used to form an overall assessment.
@@
We utilize this framework to study GPTs’ structural representation of depression for several reasons. First, depression is one of the most prevalent mental health conditions, affecting an estimated 10–15% of the population throughout their lifetime [30, 32], and is the leading cause of disability worldwide [13]. Second, there exists a vast body of research on depression’s symptoms, structure, and their inter-relationships, providing a strong theoretical reference against which to compare [55, 53, 11]. Lastly, depression is a highly heterogeneous disorder comprised of cognitive, affective, behavioral, and somatic symptoms [12]; depression is also highly co-morbid with other conditions, including anxiety [28], substance use [50], somatic disorders [19] and other internalizing conditions [21]. Thus, understanding GPT-4’s schema of depression suggests generalizations to many aspects and forms of psychopathology.
@@
While previous work on how GPTs’ assessments are formed is limited, there is growing evidence of these models’ ability to predict mental disorder [41] as compared to self-report questionnaires [14] or experts [31]. Such methods reveal what the model can do but do not explain how (i.e. explainability) it performs such tasks. The responsible use of LLMs, particularly in critical applications that involve human health and well-being, requires explaining how they perform their tasks.
@@
To address the complexities of explaining a massive model, we use a machine behavior approach [35]. Over the years, behavioral science has developed and continually refined tools and methods to explain (measure, interpret, and understand) the human mind using mathematical and empirical methods through behavioral observation [15]. Such methods have uncovered schemas — the internal concepts underlying a process — embedded within the neural system from observations of human behaviors [48, 39, 54]. Drawing parallels from cognitive and behavioral psychology, our experiments are designed to collect extensive observations of GPT-4’s ability to assess depression. We analyze its generated text (responses) and the input prompts (stimuli) to infer GPT-4’s and GPT-5’s schema of depression. This inferred schema is used to explain the models’ behavior in assessing depression.
Rahwan et al., (2019) [35] and Markowitz & Boyd (2024) [33] call for such an approach to explain GPTs’ functioning, and to the best of our knowledge, our study represents the first. We employ a robust methodological framework that can be extended to other LLMs to understand how psychopathology is represented within them. This is depicted in Figure 1 and has been described extensively under Methods.
@@
The unique contributions of this study center on explaining GPT-4’s and GPT-5’s behavior when identifying depression risk within open-ended essays. In a sample of 955 individuals, we employ measurement theory to infer GPTs’ schema of depression by analyzing (a) the relationships among key depression symptoms and (b) the role of each symptom in determining the degree of depression within the language model. This inferred schema is compared with one derived from self-reported scores to quantitatively identify similarities and differences in the conceptualization of depression as a network of interconnected symptoms. We applied differential language analysis [40, DLA] to identify the linguistic cues associated with GPT-4’s identification of depressive risk factors. Following this, we evaluated whether GPT-4 shows evidence that it uses the phrases it cited as explanations for producing its depression scores. Lastly, we used multivariate linear regression to assess how GPT-4 integrates explicitly mentioned factors in essays to estimate risks for unmentioned factors, to uncover its mechanism behind symptom estimation. We further apply this framework to GPT-5 and show that performance differences relative to GPT-4 on depression assessment are accounted for by shifts in the learned schema, linking aggregate assessment scores to interpretable internal relations among symptoms.
Figure 1: Overview of Machine Behavior Evaluation to uncover the structure of depression in GPT-4. (1) GPT-4 is prompted with step-by-step instructions on performing depression assessment on human-written essays. (2) GPT-4 carried out assessments on data from Gu et al., 2024 [16], containing depression essays written by individuals. The model estimated the severity scores for PHQ-9 symptoms explicitly mentioned in the essay, followed by scoring the remaining symptoms that were implicit in the essay. (3) GPT-4’s estimated scores and spans are used to infer its internal structure of depressive symptoms, which revealed how the web of symptoms was connected, and what language GPT-4 reliably used to estimate depression severity.
2 Methods
2.1 Data
Collection. Data comes from study participants that were recruited through Prolific[34], an online platform. From 1040 participants in total, 258 (24.8%) screened participants reported having been diagnosed with Major Depressive Disorder (MDD), 259 (24.8%) with Generalized Anxiety Disorder (GAD), and 491 (47.2%) participants were recruited from an unscreened population. 32 (3.2%) respondents were excluded for not answering the control items correctly. Removing participants that failed attention check and incomplete survey responses, the dataset comprised 963 participants, 571 identified as female, 388 as male, 2 preferred not to say, and 2 did not respond. The average age was 33.3 (SD = 11.2; range = 18 – 77) years. Participants reported being from the U.S. (n = 284) and the U.K. (n = 679).
The Swedish National Ethics Review Board deemed the research study (Ethics Application 2020-00730) exempt from requiring ethical approval according to Swedish Law (see §§ 3-4 of the Act [2003:460] on ethical review of research involving humans in Sweden). The research was carried out according to the Declaration of Helsinki. The sample comprised a mixture of screened and unscreened respondents. More details on data can be found in Gu et al.,, 2024.
Measures. Below we only describe the measures used in this study – for all details regarding the data please refer to Gu et al., 2023 [16]. Participants were asked to describe their depression using open-ended language response. The questions was adapted from previous research[23] and validated in Gu et al., 2024 [16]. Responses for the following prompt were used in this work.
Following this, participants were administered PHQ-9 [27], a 9 item questionnaire used to assess symptoms of major depressive disorder as defined in the DSM-IV [10]. The rating scale included nine items targeting symptoms such as “Feeling down, depressed, or hopeless”, referring to experiences from the last two weeks. A four-point rating scale was used (0 = Not at all to 3 = Nearly every day) to indicate the experienced severity of each symptom.
2.2 Experiment Design
Prompt. Text response describing depression was sent to GPT-4 111versioned 11/06 and referred to as gpt4-1106 in OpenAI’s GPT directory models along with the instructions to carry out assessment (made available as supplementary materials). The instruction was composed into three parts:
  1. 1.
  2. 2.
  3. 3.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
For the first two cases, the symptoms were manually corrected to match the abbreviations in Step 1 of the instructions. Type 3 generations were rectified by manually reformatting, and the summaries in type 4 were removed. The symptoms in type 5 generations were left out of the analysis.
Cleaning and Filtering GPT-4’s Responses. GPT-4 generated responses in JSON format (example output available in supplementary materials), were turned into tabular form by extracting all the scores into individual columns along with a binary variable indicating whether a symptom was found to be explicit by GPT-4. The excerpts from the user-written text identified by GPT-4 as explicit symptom markers were extracted using simple regular expressions.
GPT-5 also used the same experimental design and processing steps.
Expert Judgments. Two psychology experts independently assessed the depression severity from the depression texts on a random subsample of 209 instances. They were instructed to read the essay to identify excerpts indicative of PHQ-9 symptoms and estimate a severity score for them. If there were no indicative excerpts, they were asked to estimate the symptom severity based on the essay and the estimated explicit symptom scores. The symptoms accompanied by excerpts would pertain to explicit markers as identified by experts.
2.3 Analytic Plan
Our analytical approach draws from measurement theory and behavioral analysis to systematically understand GPT-4’s and GPT-5’s representation of depressive symptoms. We used three strategies to bring out different aspects of how the models processes and integrates depressive symptoms from natural language. We begin by assessing the convergent validity between GPT-4 and GPT-5 estimates, self-reported PHQ-9 scores, and expert judgments using Pearson correlations to establish the basic reliability of language models’ assessments before examining its underlying mechanisms. Following this, the rest of the analysis in the main paper focuses on GPT-4, and GPT-5’s analysis can be found in the Supplementary section. We perform latent structure analysis to model how symptoms are related within GPT-4. Subsequently, we use Differential Language Analysis to gain a deeper understanding of how it views depression in language, and forms severity assessments. Finally, we model the process behind the estimation of implicit symptoms’ severity based on explicit symptom estimates.
2.3.1 Latent Structure Analysis
Inter-symptom Relationships. To understand whether GPT-4’s assessment behavior reflects symptom associations consistent with established depression research, we computed Pearson correlations between all pairs of the nine PHQ-9 symptoms for both GPT-4 estimates and self-reported scores. This reveals how GPT-4 conceptualizes the relationships among different depressive symptoms compared to how individuals experience and report these relationships. The correlation matrices allow us to identify whether GPT-4 has learned symptom associations that align with or deviate from established psychometric patterns in depression research.
To contrast the inter-symptom relationships between GPT-4 and self-report, we computed the difference between the inter-symptoms relationship matrices. For statistical robustness, this difference matrix was computed over 500 trials by bootstrapped resampling to extract the 95% confidence interval of the differences. CIs that included 0 were regarded statistically insignificant difference in relationship between the two symptoms $(p>.95)$ across GPT-4 and self-report.
Differential Item Functioning Analysis. We conducted Differential Item Functioning (DIF) analysis using Item Response Theory to compare how effectively each symptom discriminates between varying levels of depression severity across GPT-4 estimates versus self-reported scores. In IRT, data from ordinal items are modeled using latent variables with logistic link functions [49]. For this analysis, we used the 1-parameter logistic IRT model (Rasch model) [36] which constrains the slope parameter to 1.0, with item location parameters estimated from responses. The item location parameter indicates where each symptom functions most effectively along the latent depression continuum. DIF analysis examines whether items function differently across groups by comparing rank orderings of item locations.
Lower ranks indicate symptoms that discriminate better at lower depression severity levels, while higher ranks indicate symptoms most informative for distinguishing individuals with severe depression. We computed DIF ranks for both GPT-4 and self-report assessments, with rank differences revealing whether GPT-4’s language-based assessment emphasizes different symptoms than traditional self-report measures across the depression severity spectrum. To ensure statistical robustness, we estimated 95% confidence intervals around rank differences using bootstrapped resampling over 500 trials, allowing us to identify statistically significant differences in how symptoms function across assessment methods.
2.3.2 Explicit Symptom Identification and Linguistic Analysis
Language Analysis of Depression Markers. Beyond examining GPT-4’s numerical estimates, we investigated whether the model’s scoring behavior aligns with linguistic patterns known to characterize depression in human language, particularly focusing on language features that extend beyond GPT-4’s explicitly identified symptom markers. For each participant’s text, we computed 1-, 2-, and 3-gram relative frequencies, categorizing tokens into two groups: those present within GPT-4’s explicit spans (text segments the model cited as evidence) and those outside these spans. To ensure robust analysis, we filtered out N-grams used by fewer than 10 participants, and removed any non alphabetic characters (including numbers or special characters). We then computed Cohen’s d [6] effect sizes comparing n-gram relative frequencies within explicit spans versus the remainder of user text. This analysis identifies linguistic features that GPT-4 systematically attributes as an explicit marker of depression.
To control for multiple comparisons, we applied Benjamini-Hochberg correction, which maintains acceptable false discovery rates when testing numerous linguistic features simultaneously. We further refined the analysis by filtering features using Pointwise Mutual Information [5] with a threshold of 4.0, which identifies n-grams with strong associative relationships to explicit symptom markers while removing spurious associations. Finally, we correlated the relative frequencies of these linguistically significant features with GPT-4’s total PHQ-9 scores to determine whether the model’s severity estimates rely on language patterns beyond its explicitly cited evidence.
2.3.3 Mechanism Analysis: Explicit-Implicit Symptom Interplay
Multivariate Regression Modeling of Symptom Inference. To understand GPT-4’s mechanism for estimating symptoms that are not explicitly mentioned in user text, we modeled how the GPT-4’s implicit symptom scores relate to its explicit symptom identifications. For each symptom $i$ from the complete set of PHQ-9 symptoms $P$, we fitted a separate multivariate linear regression model predicting GPT-4’s estimated score for symptom $i$, specifically in cases where that symptom was not explicitly identified by the model.
The regression models used scores of all other explicitly identified symptoms as predictors, creating a comprehensive picture of how GPT-4 integrates explicit information to infer unmentioned symptoms. When a predictor symptom was not explicitly identified, we replaced its score with the dataset mean for that symptom. $\beta$ represent the strength and direction of association between each explicit symptom and the implicit estimation of the target symptom:
This approach reveals GPT-4’s learned patterns about symptom co-occurrence, similar to how a clinician might infer risk for sleep problems when someone explicitly reports severe mood symptoms. To ensure statistical robustness and account for sampling variability, we repeated this entire modeling process 500 times using bootstrapped samples of the data. Coefficients whose 90% confidence intervals included zero were considered statistically non-significant and excluded from interpretation, providing conservative estimates of GPT-4’s symptom inference mechanisms.
@@ 3 Results
GPT-4 analyzed language samples describing depressive feelings from 955 individuals to estimate symptom-level scores based on the PHQ-9 questionnaire [27]. First, the model identified explicitly described symptoms, citing relevant text spans from the language sample and assigning severity scores. For the remaining PHQ-9 symptoms that were not explicitly mentioned, GPT-4 was instructed to provide rationales along with its inferred severity estimates. Participants independently completed the PHQ-9 questionnaire reflecting symptom severity over the previous two weeks. Additionally, two psychology experts manually annotated symptoms in a random subsample of 209 responses following the same guidelines given to GPT-4. GPT-4 estimates were compared against self-reported scores on the full dataset and against expert annotations in the subsample (Table 1).
@@
Table 1 shows the average Pearson correlation ($r$) between GPT-4 estimates and experts for each symptom, as well as the correlation between GPT-4 and self-reported scores, both computed on the subset assessed by the experts ($N=209$). GPT-4 demonstrated high agreement with self-report ($r=0.70$) and experts-judgments (avg $r=0.81$) on total PHQ-9 ratings. Across all symptoms, the assessors of depression essays (i.e., GPT-4 and the experts) demonstrated higher agreement with each other than either did with self-reported PHQ-9 scores. The absolute difference between GPT-4’s agreements with self-report and experts (difference between first two columns) was less than $0.10$ for all somatic symptoms (fatigue, appetite, and psychomotor agitation/retardation), except for issues related to Sleep ($\Delta=0.15$). All cognitive and emotional symptoms except for anhedonia had a difference in correlation scores over 10 Pearson points. Highest differences were seen with suicidal ideation ($\Delta=0.32$), worthlessness/guilt ($\Delta=0.20$) and concentration difficulties ($\Delta=0.14$). GPT-4’s agreement with experts ($r_{exp1}=0.80,r_{exp2}=0.81$) on total PHQ-9 score was very similar to the agreement between the experts ($r=0.83$). GPT-5 showed consistently lower agreement with self-report than GPT-4 across all symptoms and overall PHQ-9 score.
@@
The symptom-symptom associations from GPT-4 estimates in 2(a) (left) had positive Pearson correlations across the board (r=0.230.78r=0.23-0.78). This was in agreement with the symptom-symptom associations from self-reported (right) scores (r=0.380.78r=0.38-0.78). The effect sizes of the symptom-symptom relationships from GPT-4 estimates were similar to self-report for all but two symptoms, namely, psychomotor agitation/retardation and suicidal ideation. 2(b) clearly shows this difference between symptom-symptom associations from GPT-4 and self-report is statistically significant (95% CI doesn’t include 0.0). GPT-4-estimated severities of psychomotor agitation/retardation had a stronger association with other depressive symptoms than self-report. In contrast, suicidal ideation was less related to other depressive symptoms for GPT-4 than in the self-report. This is evident from the significantly lower association of suicidal ideation with the rest of the symptoms than what was observed in self-reported scores.
@@
We conducted Differential Item Functioning (DIF) analysis using Item Response Theory (IRT) to compare discriminative power between self-report and GPT-4-derived symptom scores [47, 38, 2]. DIF ranks reflect how effectively each symptom distinguishes between varying levels of latent depression severity: lower ranks indicate better discrimination at lower severity, and higher ranks indicate better discrimination at higher severity. A positive rank difference (GPT-4 minus self-report) implies that GPT-4’s language-based assessment discriminates better at higher depression severity levels compared to self-report, while a negative difference indicates the opposite. DIF ranks for both methods, along with their 95% confidence intervals estimated via bootstrapped resampling over 500 trials, are presented in 2(c).
@@
Overall, the discriminative characteristics of items were consistent between GPT-4 and self-report assessments. Suicidal ideation, appetite, and psychomotor agitation/retardation had higher ranks (7–9), indicating these symptoms were most informative in distinguishing individuals with higher latent depression severity. On the other hand, anhedonia, depressed mood, and fatigue had lower ranks (1–3), signifying these symptoms effectively discriminated between individuals experiencing milder or lower levels of depression. All item ranks showed narrow 95% confidence intervals, reflecting high reliability. Suicidal ideation (rank = 9), sleep (rank = 4), and anhedonia (rank = 3) exhibited identical ranks between GPT-4 and self-report. Rank differences for other symptoms were minor, reinforcing that language-based severity scores from GPT-4 closely align with the discriminative item characteristics observed in self-reported PHQ-9 scores.
Explicit vs. Implicit Symptoms: GPT-4 Estimates Explicitly Mentioned Symptoms More Precisely.
@@
Agreement between GPT-4 estimates and self-report was consistently higher when symptoms were explicitly mentioned, showing an average increase of +0.18+0.18 Pearson correlation points over implicitly inferred symptoms –— even though explicit mentions were identified in only about one-quarter of instances (3(a)). Notably, explicit mentions significantly improved GPT-4’s accuracy ($\geq 0.25$ Pearson points) in estimating anhedonia, depressed mood, and suicidal ideation compared to implicit inferences. Language explicitly describing these three symptoms enabled more precise estimates from GPT-4, despite substantial variation in how frequently they were identified to be explicitly mentioned (range: 5.55%–96.44%). The cardinal symptoms, depressed mood and anhedonia, had particularly high rates of explicit mentions (96% and 53%, respectively), considerably surpassing other symptoms.
Uncovering GPT-4’s language of depression: GPT-4’s cited evidences were reliable
The left bar chart of 3(b) shows that words identified by GPT-4 as explicit markers of symptoms (positive Cohen’s $d$ values). These include emotion words (e.g., ”depressed,” ”sad,” ”down”), verbs (e.g., ”do”), and somatic descriptors (e.g., ”tired,” ”sleep,” ”energy”). In contrast, words that GPT-4 rarely cited as explicit markers (negative Cohen’s $d$) were mainly neutral functional words and nouns (e.g., ”pandemic,” ”home”). All these highlighted words showed statistically significant associations after correcting for multiple comparisons (Benjamini-Hochberg corrected, $p<0.05$).
The bar chart on the right further shows that 10 of the 17 explicitly cited words significantly correlated with GPT-4’s estimated total PHQ-9 scores ($p<0.05$). Phrases such as ”have been depressed,” ”sleep,” and ”low” (in ”low energy”) positively correlated with higher severity scores, while phrases like ”not been depressed” and ”happy” negatively correlated, indicating lower severity. Among words not explicitly cited, only ”it” showed a small but significant correlation ($r=0.09$), suggesting GPT-4 primarily bases its severity scores on explicitly cited meaningful linguistic markers.
Interplay between explicitly mentioned symptoms and inferred symptoms.